# An Embedding Theorem for tractor bundles, and an application in conformal pseudo-Riemannian geometry

### v1

@@ 1.1. Parabolic geometries and tractor bundles
A semi-Riemannian metric $g$ on $M^{n},n\geq 3$, is Einstein if Ric $\mbox{Ric }g=\frac{S}{2}g$, where Ric is the Ricci curvature tensor and $S$ is the scalar curvature. It is conformally almost-Einstein if there exists a smooth function $\sigma$, not constant 00, and a smooth function $\rho$ such that
@@
is the Schouten tensor; in other words, $\nabla^{2}\sigma+\sigma\cdot P$ is pure trace. When $\sigma$ is nonvanishing, then $\sigma^{-2}\cdot g$ is an Einstein metric in the conformal class of $g$ (see, eg, [10, Prop 3.6]).
@@
1.1.2. Projective structures.
The projective class of a torsion-free connection $\nabla$ comprises those torsion-free connections determining the same geodesic curves, up to reparamerization, and is denoted []delimited-[][$\nabla$][ ∇ ]. It corresponds to a Cartan geometry modeled on ${\bf RP}^{n}=G/P$ for $G=\mbox{SL}(n+1,{\bf R})$ and $P$ the maximal parabolic subgroup stabilizing a line in ${\bf R}^{n+1}$, for $n=\mbox{dim }M\geq 2$. The Cartan bundle is again a principal $P$-subbundle of the second-order frame bundle. The Cartan curvature vanishes if and only if $M$ admits a $G$-atlas to ${\bf RP}^{n}$—a projective structure in the classical sense.
Projective structures with nontrivial curvature are in fact also a classical topic, when they comprise the projective class of the Levi-Civita connection of a Riemannian metric. These metrizable projective structures were studied, particularly in low dimensions, in the nineteenth centry by U. Dini, E. Beltrami, T. Levi-Civita [19], and many others since then. More recently, authors such as A.S. Solodovnikov, V. Matveev, V. Kiosak, and others have studied the PDEs corresponding to projective mobility for Riemannian or pseudo-Riemannian metrics and proved rigidity results on the projective automorphisms of such structures [33, 20, 16, 5].
The metrizability of a given projective structure $(M^{n},[\nabla])$ has been found by M. Eastwood and V. Matveev [12] to correspond to the existence of parallel sections for a connection $\nabla^{\mathcal{V}}$ on the tractor bundle $\mathcal{V}=\hat{M}\times_{P}\bf V$, associated to the $G$-representation ${\bf V}=\mbox{Sym}^{2}{\bf R}^{n+1}$; it is the prolongation connection for the corresponding first BGG operator.
1.1.3. Infinitesimal automorphisms.
Let $(M,\hat{M},\omega)$ be any Cartan geometry modeled on $(\mathfrak{g},P)$. The adjoint tractor bundle is the associated vector bundle $\mathcal{A}$ for ${\bf V}=\mathfrak{g}$ with the adjoint representation. A section of $\mathcal{A}$ gives rise via the Cartan connection to a vector field on $\hat{M}$ which is invariant under the right-$P$-action, and thus descends to a vector field on $M$. A. Čap observed that for a certain connection on $\mathcal{A}$, again related to a first BGG operator, parallel sections correspond to infinitesimal automorphisms of the Cartan geometry—that is, vector fields for which the flow preserves the Cartan connection [36]. These descend to vector fields on $M$ for which the flow preserves the geometric structure encoded by the Cartan geometry, called infinitesimal automorphisms or Killing fields of the structure.
Infinitesimal automorphisms thus belong to the framework of parallel sections of natural connections on tractor bundles. The purpose of this article is to prove a generalization of the Gromov-Zimmer Embedding Theorem, which originally concerns infinitesimal automorphisms, to general parallel sections of automorphisms-invariant connections on tractor bundles. We briefly recall the Embedding Theorem for infinitesimal automorphisms.
@@ 1.2. Gromov-Zimmer Embedding Theorem
Zimmer’s Embedding Theorem first appeared in [39] in the context of a noncompact simple Lie group $H$ acting by automorphisms of a $G$-structure $\omega$ of algebraic type on a compact manifold $M^{n}$, such that $G$ also defines a volume form. In this context, it gives a linear injection $\iota:\mathfrak{h}\rightarrow{\bf R}^{n}$ and an algebraic group $\check{H}<G$ locally isomorphic to $H$ such that $\iota$ intertwines the adjoint representation $\mbox{Ad}_{\mathfrak{h}}H$ with the representation of $\check{H}$ on $\iota(\mathfrak{h})\subseteq{\bf R}^{n}$. Zimmer provided an impressive application to the case $\omega$ is a Lorentzian metric, proving that $H$ must be locally isomorphic to $\mbox{SL}(2,{\bf R})$.
M. Gromov subsequently gave a very general version of the Embedding Theorem in [14, 5.2.A], which applies to a Lie algebra $\mathfrak{h}$ of infinitesimal automorphisms of a Gromov-rigid geometric structure $\omega$ of algebraic type on a compact manifold $M$, together with a group $B\leq{\textrm{Aut}}(M,\omega)$ preserving a finite measure $\mu$ on $M$ and normalizing $\mathfrak{h}$. Combined with Gromov’s Frobenius Theorem [14, 1.3.A], the Embedding Theorem led to the Gromov-Zimmer Centralizer and Representation Theorems (see [41]).
Our version of the Embedding Theorem will be a direct generalization of the version proved by Bader, Frances, and the first author in [3, Thm 1.2] for Cartan geometries. If $(M,\hat{M},\omega)$ is a Cartan geometry modeled on $(\mathfrak{g},P)$ (see Definition 2.1 below), and $\mathfrak{h}$ is an algebra of infinitesimal automorphisms, then $\mathfrak{h}$ acts freely on $\hat{M}$, and $\iota_{\hat{x}}:X\mapsto\omega_{\hat{x}}(X_{\hat{x}})$ is a linear injection $\mathfrak{h}\rightarrow\mathfrak{g}$ for all $\hat{x}\in\hat{M}$. This version of the Embedding Theorem says that for $\mu$-almost-every $x\in M$, for all $\hat{x}\in\hat{M}_{x}$, the injection $\iota_{\hat{x}}$ intertwines the $B$-representation on $\mathfrak{h}$ with that of an algebraic subgroup $\check{B}<P$ on $\iota_{\hat{x}}(\mathfrak{h})\subseteq\mathfrak{g}$. Here $B$ is as in the previous paragraph, and some technical details have been omitted, to be included in the next subsection.
@@ 1.3. Results
As an application, we prove the following rigidity theorem in conformal pseudo-Riemannian geometry, partially extending a previous result of Pecastaing for conformal Lorentzian actions to higher signatures.
@@
This result was proved for $p^{\prime}=1$ without assuming $(M,g)$ simply connected or real-analytic in [28, Cor 1]. For our proof we apply the Embedding Theorem to two tractor bundles simultaneously: to the conformal Killing fields, as in the existing version of the theorem, and also to the conformal-to-almost-Einstein scales, corresponding to the parallel sections of the standard conformal tractor bundle.
From the embedding in the first bundle, we prove Proposition 4.12 to obtain a special 1-parameter subgroup in $\check{B}$, which is of balanced linear type. Then we apply the Frobenius Theorem to ensure that this 1-parameter subgroup is the isotropy of a 1-parameter group of local conformal transformations fixing $x$. Such a conformal flow in turn implies the existence of local tractor solutions, more precisely of a 2-parameter family of Ricci-flat metrics in the conformal class restricted to a neighborhood of $x$. Analyticity of the metric and $1$-connectedness of $M$ allow us to extend these solutions to all of $M$, even though none of the Ricci-flat metrics may extend to a global metric on $M$.
@@ 2.1.1. Cartan geometries and parabolic geometries
A Cartan geometry of type $({\mathfrak{g}},P)$ on a manifold $M$ is given by a principal $P$-bundle $\pi:\hat{M}\rightarrow M$ equipped with a Cartan connection, that is, a $P$-equivariant $1$-form $\omega\in\Omega^{1}(\hat{M},{\mathfrak{g}})$ such that
@@
The homogeneous model of a Cartan geometry as in Definition 2.1 is the homogeneous space $G/P$ equipped with the natural projection $G\rightarrow G/P$ and the Maurer–Cartan form $\omega_{G}\in\Omega^{1}(G,{\mathfrak{g}})$ of $G$, which encodes the natural trivialization $TG\cong G\times{\mathfrak{g}}$ by left-invariant vector fields. The curvature of a Cartan geometry is given by the following semibasic, ${\mathfrak{g}}$-valued, $P$-equivariant $2$-form on $\hat{M}$:
$K(\xi,\eta)=d\omega(\xi,\eta)+[\omega(\xi),\omega(\eta)]\quad\textrm{ for }\xi,\eta\in\mathfrak{X}(\hat{M}).$
@@
Sections of the adjoint bundle $\mathcal{A}=\hat{M}\times_{P}\mathfrak{g}$ from Section 1.1.3 correspond via $\omega$ to right-$P$-invariant vector fields on $\hat{M}$, which descend to vector fields on $M$. The Cartan connection $\omega$ induces isomorphisms of vector bundles
@@
These identifications can easily be derived from the axioms for $\omega$ (see also [32, Thm 5.3.15]). Because the curvature $K$ is semibasic and $P$-equivariant, it descends to a $2$-from on $M$ with values in $\mathcal{A}$, which, via (2), can be identified with a $P$-equivariant Cartan curvature function
@@
Many important geometric structures admit equivalent descriptions as Cartan geometries, in particular semi-Riemannian metrics. A broad and interesting class of examples are the geometric structures underlying parabolic geometries, which are Cartan geometries of type $({\mathfrak{g}},P)$, where $P<G$ is a parabolic subgroup of a semisimple Lie group. These correspond to certain bracket-generating distributions on $M$ endowed with a geometric structure defined by $G_{0}$, the reductive complement in a Levi decomposition of $P\cong G_{0}\ltimes P_{+}$. These structures correspond canonically to normal parabolic Cartan geometries. See [8, 24, 34] for the precise general statements, including the notions of normal and regular. In the table below we list some of the most prominent examples of geometric structures admitting an equivalent description as regular normal parabolic geometries of type $({\mathfrak{g}},P)$. For more examples, see [9]. A concise reference on the construction of the Cartan connection in conformal or projective geometry is [17, Ch IV]. All parabolic geometries will be assumed to be regular and normal throughout the sequel.
@@
Almost quaternionic $4n$-manifolds $\mathfrak{sl}(n+1,\mathbb{H})$ stabilizer of a quaternionic
@@
Generalized path geometries $\mathfrak{sl}(n+2,{\mathbb{R}})$ stabilizer in ${\bf R}^{n+2}$ of a flag
on $12n+$ 1-manifolds, $n\geq 2$ of line inside a plane
@@ 2.1.2. Tractor connections
Viewing $t$ in $C^{\infty}(\hat{M},{\bf W})^{P}$ and $s$ in $C^{\infty}(\hat{M},{\mathfrak{g}})^{P}$, the section $D_{s}t\in C^{\infty}(\hat{M},{\bf W})^{P}$ evaluates at $\hat{x}\in\hat{M}$ to the derivative of $t$ in the direction of $\omega_{\hat{x}}^{-1}(s(\hat{x}))$. The tractor connection on a tractor bundle $\mathcal{V}$ can then be expressed in terms of the fundamental derivative as follows:
@@ 2.1.5. The conformal tractor connection
Let $(M,g)$ be a semi-Riemannian manifold of dimension $n\geq 3$ and let $(M,\hat{M},\omega)$ and $\mathcal{V}$ be the corresponding parabolic Cartan geometry with model $({\mathfrak{g}},P)$ and conformal tractor bundle, respectively, as in Section 1.1.1. Recall that $P$ is the stabilizer of an isotropic line ${\bf V}_{1}\subset{\bf V}\cong{\bf R}^{p+1,q+1}$. Setting ${\bf V}^{1}={\bf V}_{1}$ and ${\bf V}^{0}={\bf V}_{1}^{\perp}$, there is a $P$-invariant filtration ${\bf V}^{1}\subset{\bf V}^{0}\subset{\bf V}^{-1}={\bf V}$, cf. (3), giving rise to an intrinsic filtration $\mathcal{V}^{1}\subset\mathcal{V}^{0}\subset\mathcal{V}$.
@@
in which $\mathcal{V}_{0}\cong TM$ and $\mathcal{V}_{-1}$ is identified with the line bundle of conformal scales. References for the formulas in this section are [10, Eqns 22, 32]. With respect to a decomposition (7), the standard conformal tractor connection $\nabla=\nabla^{\omega}$ is
@@
grad $\nabla\mbox{grad }\sigma+\sigma P^{\sharp}=-\rho\cdot\mbox{Id}$
@@ 2.2. Borel Density Theorem
We use a version of the Borel Density Theorem which can be found in [3, Cor 3.2], and which is due in this generality to Y. Shalom [31, Thm 3.11] (see also [11, Thm 2.6]). The setting is a locally compact group $S$ acting on an algebraic variety $\mathbb{W}$ via a continuous homomorphism $\rho:S\rightarrow\mbox{Aut }\mathbb{W}$. We will denote by $\bar{S}$ the Zariski closure of $\rho(S)$ in Aut $\mbox{Aut }\mathbb{W}$.

### v2

@@ 1.1. Parabolic geometries and tractor bundles
A semi-Riemannian metric $g$ on $M^{n},n\geq 3$, is Einstein if $\mbox{Ric }g=\frac{S}{2}g$, where Ric is the Ricci curvature tensor and $S$ is the scalar curvature. It is conformally almost-Einstein if there exists a smooth function $\sigma$, not constant 0, and a smooth function $\rho$ such that
@@
is the Schouten tensor; in other words, $\nabla^{2}\sigma+\sigma\cdot P$ is pure trace. When $\sigma$ is nonvanishing, then $\sigma^{-2}\cdot g$ is an Einstein metric in the conformal class of $g$ (see, eg, [11, Prop 3.6]).
@@
There are many other significant geometric PDEs on conformal manifolds, and more generally on manifolds equipped with parabolic geometries, whose solutions prolong to parallel sections of the prolongation connection of a first BGG-operator; on conformal manifolds this applies for instance also to the equations for conformal Killing forms [16, 31] and for twistor spinors [5], and for projective and c-projective structures to the equation controlling the metrizability of these structure [13, 7], among many others.
Infinitesimal automorphisms of geometric structures underlying Cartan geometries also belong to the framework of parallel sections of natural connections on tractor bundles. Let $(M,\hat{M},\omega)$ be any Cartan geometry modeled on $(\mathfrak{g},P)$. The adjoint tractor bundle is the associated vector bundle $\mathcal{A}$ for ${\bf V}=\mathfrak{g}$ with the adjoint representation. A section of $\mathcal{A}$ gives rise via the Cartan connection to a vector field on $\hat{M}$ which is invariant under the right-$P$-action, and thus descends to a vector field on $M$. A. Čap observed that for a certain connection on $\mathcal{A}$, again related to an intrinsically defined differential operator on $M$, parallel sections correspond to infinitesimal automorphisms of the Cartan geometry—that is, vector fields for which the flow preserves the Cartan connection [36]. These descend to vector fields on $M$ for which the flow preserves the geometric structure encoded by the Cartan geometry, called infinitesimal automorphisms or Killing fields of the structure.
This article concerns a generalization of the Gromov-Zimmer Embedding Theorem, which originally pertains to infinitesimal automorphisms, to general parallel sections of automorphism-invariant connections on tractor bundles. We briefly recall the Embedding Theorem for infinitesimal automorphisms.
@@ 1.2. Gromov-Zimmer Embedding Theorem
Zimmer’s Embedding Theorem first appeared in [39] in the context of a noncompact simple Lie group $H$ acting by automorphisms of a $G$-structure $\omega$ of algebraic type on a compact manifold $M^{n}$, such that $G$ also defines a volume form. In this context, the theorem gives a linear injection $\iota:\mathfrak{h}\rightarrow{\bf R}^{n}$ and an algebraic group $\check{H}<G$ locally isomorphic to $H$ such that $\iota$ intertwines the adjoint representation $\mbox{Ad}_{\mathfrak{h}}H$ with the representation of $\check{H}$ on $\iota(\mathfrak{h})\subseteq{\bf R}^{n}$; in particular, $H$ locally embeds in the structure group $G$. Zimmer provided an impressive application for $\omega$ a Lorentzian metric, proving that $H$ must be locally isomorphic to $\mbox{SL}(2,{\bf R})$.
M. Gromov subsequently gave a very general version of the Embedding Theorem in [17, 5.2.A], which applies to a Lie algebra $\mathfrak{h}$ of infinitesimal automorphisms of a Gromov-rigid geometric structure $\omega$ of algebraic type on a compact manifold $M$, together with a group $B\leq{\textrm{Aut}}(M,\omega)$ preserving a finite measure $\mu$ on $M$ and normalizing $\mathfrak{h}$. Combined with Gromov’s Frobenius Theorem [17, 1.3.A], the Embedding Theorem led to the Gromov-Zimmer Centralizer and Representation Theorems (see [41]).
[…]
@@ 1.3. Results
As an application, we prove the following rigidity theorem in conformal pseudo-Riemannian geometry.
@@
Without assuming $(M,g)$ simply connected or real-analytic, V. Pecastaing proved this result for $p^{\prime}=1$ [28, Cor 1]. He proved in [29, Thm 3] that any simple group $H<\mbox{Conf}(M,g)$, for $M$ closed and $g$ Lorentzian, locally embeds in $\mbox{O}(2,n+1)$. In analogy with Zimmer’s result on volume-preserving actions of simple groups in Section 1.2 above, the following question arises:
Question 1.1.
Let $(M,\omega)$ be a closed manifold with a parabolic geometry modeled on $(\mathfrak{g},P)$. Suppose that $H\leq{\textrm{Aut}}(M,\omega)$ is a simple Lie group. Does the Lie algebra $\mathfrak{h}$ of $H$ embed in $\mathfrak{g}$?
For the proof of Theorem 1.2 we apply the Embedding Theorem to two tractor bundles simultaneously: to the conformal Killing fields, as in the usual version of the theorem, and also to the conformal-to-almost-Einstein scales, corresponding to the parallel sections of the standard conformal tractor bundle. From the first embedding, we prove the algebraic Proposition 4.12, which says $p\geq 2p^{\prime}-1$; in the case of equality, it gives $p<q$ as well as precise information about how the positive root spaces of $\mathfrak{h}$ embed in those of $\mathfrak{g}$. Proposition 4.12 demonstrates how in an extremal case, and with $H$ simple, careful algebraic analysis of the output of the classical Embedding Theorem yields that it is very close to a Lie algebra embedding of $\mathfrak{h}$ into $\mathfrak{g}$.
@@
If $H$ locally isomorphic to $SU(p^{\prime},q^{\prime})$ acts isometrically on $(M,g)$ as above, then by Zimmer’s result there is a Lie algebra embedding $\mathfrak{su}(p^{\prime},q^{\prime})\hookrightarrow\mathfrak{so}(p,q)$, which implies $p\geq 2p^{\prime}$ (see Proposition 4.12). Thus the extremal case $p=2p^{\prime}-1$ of Theorem 1.2 corresponds to essential conformal transformations. A version of Question 1.1 for conformal transformations, in particular for essential conformal groups, is discussed in [29]. See [4], [15] for important results in this direction in terms of the real rank of $H$.
@@ 2.1.1. Cartan geometries and parabolic geometries
A Cartan geometry modeled on $({\mathfrak{g}},P)$ on a manifold $M$ is given by a principal $P$-bundle $\pi:\hat{M}\rightarrow M$ equipped with a Cartan connection, that is, a $P$-equivariant 11-form $\omega\in\Omega^{1}(\hat{M},{\mathfrak{g}})$ such that
@@
The homogeneous model of a Cartan geometry as in Definition 2.1 is the homogeneous space $G/P$ equipped with the natural projection $G\rightarrow G/P$ and the Maurer–Cartan form $\omega_{G}\in\Omega^{1}(G,{\mathfrak{g}})$ of $G$, which encodes the natural trivialization $TG\cong G\times{\mathfrak{g}}$ by left-invariant vector fields. The curvature of a Cartan geometry is given by the following semibasic, $\mathfrak{g}$}-valued, $P$-equivariant 22-form on $\hat{M}$:
[…]
@@
Sections of the adjoint bundle $\mathcal{A}=\hat{M}\times_{P}\mathfrak{g}$ from Section 1.1 correspond via $\omega$ to right-$P$-invariant vector fields on $\hat{M}$, which descend to vector fields on $M$. The Cartan connection $\omega$ induces isomorphisms of vector bundles
@@
These identifications can easily be derived from the axioms for $\omega$ (see also [33, Thm 5.3.15]). Because the curvature $K$ is semibasic and $P$-equivariant, it descends to a $2$-from on $M$ with values in $\mathcal{A}$, which, via (2), can be identified with a $P$-equivariant Cartan curvature function
@@
Many important geometric structures admit equivalent descriptions as Cartan geometries, in particular semi-Riemannian metrics. A broad and interesting class of examples are the geometric structures underlying parabolic geometries, which are Cartan geometries of type $({\mathfrak{g}},P)$, where $P<G$ is a parabolic subgroup of a semisimple Lie group. These correspond to certain bracket-generating distributions on $M$ endowed with a geometric structure defined by $G_{0}$, the reductive complement in a Levi decomposition of $P\cong G_{0}\ltimes P_{+}$. These structures correspond canonically to normal parabolic Cartan geometries. See [9, 24, 34] for the precise general statements, including the notions of normal and regular. In the table below we list some of the most prominent examples of geometric structures admitting an equivalent description as regular normal parabolic geometries of type $({\mathfrak{g}},P)$. For more examples, see [10]. A concise reference on the construction of the Cartan connection in conformal or projective geometry is [19, Ch IV]. All parabolic geometries will be assumed to be regular and normal throughout the sequel.
@@
Almost quaternionic $4n$-manifolds $\mathfrak{sl}(n+1,{\bf H})$ stabilizer of a quaternionic
@@
Generalized path geometries $\mathfrak{sl}(n+2,{\bf R})$ stabilizer in ${\bf R}^{n+2}$ of a flag
on $2n+1$-manifolds, $n\geq 2$ of line inside a plane
@@ 2.1.2. Tractor connections
Viewing $t$ in $C^{\infty}(\hat{M},{\bf W})^{P}$ and $s$ in $C^{\infty}(\hat{M},{\mathfrak{g}})^{P}$, the section $D_{s}t\in C^{\infty}(\hat{M},{\bf W})^{P}$ evaluates at $\hat{x}\in\hat{M}$ to the derivative of $t$ in the direction of $)\omega_{\hat{x}}^{-1}(s(\hat{x}$)). The tractor connection on a tractor bundle $\mathcal{V}$ can then be expressed in terms of the fundamental derivative as follows:
@@ 2.1.5. The conformal tractor connection
Let $(M,g)$ be a semi-Riemannian manifold of dimension $n\geq 3$ and let $(M,\hat{M},\omega)$ and $\mathcal{V}$ be the corresponding parabolic Cartan geometry with model $({\mathfrak{g}},P)$ and conformal tractor bundle, respectively, as in Section 1.1. Recall that $P$ is the stabilizer of an isotropic line ${\bf V}_{1}\subset{\bf V}\cong{\bf R}^{p+1,q+1}$. Setting ${\bf V}^{1}={\bf V}_{1}$ and ${\bf V}^{0}={\bf V}_{1}^{\perp}$, there is a $P$-invariant filtration ${\bf V}^{1}\subset{\bf V}^{0}\subset{\bf V}^{-1}={\bf V}$, cf. (3), giving rise to an intrinsic filtration $\mathcal{V}^{1}\subset\mathcal{V}^{0}\subset\mathcal{V}$.
@@
in which $\mathcal{V}_{0}\cong TM$ and $\mathcal{V}_{-1}$ is identified with the line bundle of conformal scales. References for the formulas in this section are [11, Eqns 22, 32]. With respect to a decomposition (7), the standard conformal tractor connection $\nabla=\nabla^{\omega}$ is
@@
$\nabla\mbox{grad }\sigma+\sigma P^{\sharp}=-\rho\cdot\mbox{Id}$
@@ 2.2. Borel Density Theorem
We use a version of the Borel Density Theorem which can be found in [3, Cor 3.2], and which is due in this generality to Y. Shalom [32, Thm 3.11] (see also [12, Thm 2.6]). The setting is a locally compact group $S$ acting on an algebraic variety $\mathbb{W}$ via a continuous homomorphism $\rho:S\rightarrow\mbox{Aut }\mathbb{W}$. We will denote by $\bar{S}$ the Zariski closure of $\rho(S)$ in $\mbox{Aut }\mathbb{W}$.